a businessman invested Php. 16,000.000, part at 9% and the rest at 4%. His total interest from both investments was Php. 990.00. How much did he invest at each rate?

Answer by mananth(16946)

Investment I 9.00% per annum ---x
Investment II 4.00% per annum ---y

x + y= 16000000 ------------------------1
9.00% x + 4.00% y = 990000
Multiply by 100
9 x + 4 y = 99000000--------2

Multiply (1) by -9
we get

-9 x -9 y = -144000000.00

Add this to (2)

0 x -5 y = -45000000

divide by -5

y = 9000000 investment at 4.00%
Balance 7000000 investment at 9.00%

CHECK
7000000 630000.00
9000000 360000.00
Total 990000.00